One angle of a right triangle measures 75 degrees, what is the measure of the other acute angle?
Found 2 solutions by solver91311, edjones:
Answer by solver91311(12126)

The total of the measures of the three angles of ANY triangle is 180 degrees. You are given a right triangle, so one of the three angles measures 90 degrees by definition, and you are given that another one of the angles measures 75 degrees. So add 90 to 75 and subtract the sum from 180 to get the measure of the third angle. You get to do your own arithmetic.

Answer by edjones(7311)

90+75+x=180
165+x=180
x=180-165
x=15
.
Ed